OpenStudy (anonymous):

Help with this math problem please! :c (MEDALS) What is the greatest value in the range of f(x) = x^2 - 3 for the domain { -3, 0, 1, 2}, please explain how you got the answer

OpenStudy (anonymous):

\[f(-3)=(-3)^2-3=9-3=6\] repeat for the other three

OpenStudy (anonymous):

So it would be 6 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Great, thank you